Circular No. 7216

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999cw IN MCG –01-02-001 L. Rizzi, Universitá di Padova; F. Patat, European Southern Obser-

vatory (ESO); S. Benetti, Telescopio Nazionale Galileo; and E. Cappellaro and M. Turatto, Osservatorio Astronomico di Padova, write: “V. Ripepi and M. Marconi (Osservatorio di Capodimonte) have observed SN 1999cw with the Danish 1.5-m telescope (+ DFOSC; 17-min exposure, range 500– 970 nm, resolution 1.0 nm) on July 1.39 UT at La Silla. A full reduction of the CCD spectrogram shows this object (IAUC 7211) to be a type-Ia supernova around maximum light. The spectrum is dominated by P-Cyg profiles superimposed on a blue continuum, with the Si II 635.5-nm line particularly weak, similar to the peculiar type-Ia SN 1991T. The expansion velocity deduced from the minimum of the Si II line is 11 000 km/s, which is consistent with SN 1991T at a similar epoch. The parent galaxy’s recession velocity is 3725 ± 39 km/s (Huchra et al. 1993, A.J. 105, 1637).”

SUPERNOVA 1999cz IN NGC 5078 Rizzi et al. also report: “E. Pompei has observed SN 1999cz with the

ESO-Danish 1.54-m telescope (+ DFOSC) on July 6.08 UT at La Silla. A full reduction of the CCD spectrogram (range 400–900 nm, resolution 1.0 nm) shows that this object (IAUC 7214) is indeed a supernova. Weak Si II (635.5 nm), very-well-developed P-Cyg profiles, and a rather red continuum are consistent with a type-Ic supernova, 3–4 weeks after maximum light.

The spectrum (http://sc6.sc.eso.org/~lrizzi/sn99cz spec.gif) is

similar to that of SN 1997B (ESO database, unpublished) at 1 month past maximum, even if the near-infrared Ca triplet is much fainter in SN 1999cz. The most prominent features are identified as Si II, Ca II, Na I D, and Fe II. The expansion velocities, deduced from the minima of Si II 635.5-nm and Na I D, are 6900 and 9200 km/s, respectively. The spectrum appears to be heavily reddened. A very strong, unresolved absorption feature (EW = 0.58 nm) is visible at the parent galaxy rest-frame wavelength of Na I D. This is consistent with the position of the supernova, which appears to be projected just above the dust lane of NGC 5078.”

V382 VELORUM Photometry of V382 Vel by P. M. Kilmartin at Mount John (reference

star HD 92518; cf. IAUC 7196, 7209): June 26.341 UT, V = 7.30, U − B = −0.80, B − V = −0.14, V − R = +1.46, V − I = +0.90; 28.351, 7.35, −0.94, −0.14, +1.44, +0.86; July 4.382, 7.46, −0.94, −0.14, +1.41, +0.77.

1999 July 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green